# Torsion of Rational Elliptic Curves over the Cyclotomic Extensions of $\mathbb{Q}$

### v5

@@ 4. Isogenies
  • If $C_{3}(E)=4$, then there exists a representative of the class of $E$ with a $\mathbb{Q}$-rational cyclic subgroup of order $27$, and $C_{2}(E)=1$ so $C(E)\leq 4$.
@@
This corollary is also used as a consequence of the Néron–Ogg–Shafarevich criterion in the literature.
@@
Assume that $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/\ell\mathbb{Z}$ for some prime $p$ and some $\ell\in\{11,19,37,43,67,163\}$. By Lemma 4.3, $E$ must have an $\ell$-isogeny. In Chapter 7 of [3], the $j$-invariants of such elliptic curves are listed. In total, there are $7$ different $j$-invariants, and we can also identify the Cremona labels of the corresponding elliptic curves. Using the LMFDB database, we observe that the conductors of these elliptic curves are equal to $N_{E}=\ell^{2}$ when working with the minimal quadratic twists of the corresponding $j$-invariants. This fact, combined with Corollary 4.6, implies that $\mathbb{Q}(E[\ell])$ is unramified at all primes other than $\ell$. For the remainder of the proof, the elliptic curve $E/\mathbb{Q}$ will always refer to the minimal quadratic twist of its $j$-invariant.
@@ 5. Relaxing the Conditions
In this section, we will discuss the implications of relaxing the conditions on our prime $p$ when classifying $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$. In our first theorem, we classified all possible torsion subgroups when $p-1$ is not divisible by $3$, $4$, or $5$. Now, we will gradually relax these conditions. As a result, the number of groups appearing in the classification will increase. One of the first theorems we obtain is the following.
@@
Let $E/\mathbb{Q}$ be an elliptic curve. Let $p>3$ be a prime such that $p-1$ is not divisible by $3$ or $4$. Then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is either one of the groups from Mazur’s Theorem, or one of the following groups:
/2×/10,/2×/12,/16,/25,or $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z},\;\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z},\;\mathbb{Z}/16\mathbb{Z},\;\mathbb{Z}/25\mathbb{Z},\;\text{or }\;\mathbb{Z}/11\mathbb{Z}.$
Moreover, the group $\mathbb{Z}/11\mathbb{Z}$ is realized only when $p=11$.
@@
The proof is the same as that of Theorem 1.1. The only difference is in the application of Lemma 3.2 in the proof. Since we have removed the condition that $p-1$ is not divisible by $5$, we cannot conclude that $\mathbb{Z}/11\mathbb{Z}$ or $\mathbb{Z}/25\mathbb{Z}$ are not realizable. We can use Lemma 4.7 to see that $11$-torsion is realized over $\mathbb{Q}(\zeta_{p})$ only when $p=11$ and only for $3$ different $j$-invariants (see Table 1). The elliptic curve 11.a3 satisfies $E(\mathbb{Q}(\zeta_{11}))_{\text{tors}}\cong\mathbb{Z}/25\mathbb{Z}$. We could not find any other prime $p$ satisfying $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/25\mathbb{Z}$ for some elliptic curve $E/\mathbb{Q}$, and it is beyond our scope to show that no other primes exist. ∎
@@
Let $E/\mathbb{Q}$ be an elliptic curve. Let $p>3$ be a prime such that $p-1$ is not divisible by $3$. Then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is either one of the groups from Mazur’s Theorem, or one of the following groups:
$E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\simeq\begin{cases}\mathbb{Z}/N\mathbb{Z}&\text{ with }1\leq N\leq 13,\text{ or }N=15,16,25,\\ \mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}&\text{ with }1\leq N\leq 6,\text{ or }N=8.\\ \end{cases}$
Moreover, the group $\mathbb{Z}/11\mathbb{Z}$ is realized only when $p=11$.
@@
The proof is the same as in Theorem 1.1. The only difference lies in the use of Lemma 3.2 during the proof. Since we removed the condition that $p-1$ is not divisible by $4$ or $5$, we cannot rule out the possibility that $\mathbb{Z}/N\mathbb{Z}$ is realizable for $N=11,13,15,$ or $25$. Similarly, we cannot exclude the possibility of $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$. We also cannot rule out the cases $\mathbb{Z}/17\mathbb{Z}$ and $\mathbb{Z}/37\mathbb{Z}$ using Lemma 3.2, but Lemma 4.8 allows us to exclude them.
By Lemma 4.7, $11$-torsion is realized over $\mathbb{Q}(\zeta_{p})$ only when $p=11$. The elliptic curve 50.b3 satisfies $E(\mathbb{Q}(\zeta_{5}))_{\text{tors}}\cong\mathbb{Z}/15\mathbb{Z}$. The elliptic curve 11.a3 satisfies $E(\mathbb{Q}(\zeta_{11}))_{\text{tors}}\cong\mathbb{Z}/25\mathbb{Z}$.
We have no examples of $\mathbb{Z}/13\mathbb{Z}$ when $p\equiv 2\pmod{3}$, or of $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$ for any prime $p$, and our methods are not effective in ruling out their existence. ∎
@@
Let $E/\mathbb{Q}$ be an elliptic curve. Let $p>7$ be a prime. Then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is either one of the groups from Mazur’s Theorem, or one of the following groups:
$E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\simeq\begin{cases}\mathbb{Z}/N\mathbb{Z}&\text{ with }1\leq N\leq 13,\text{ or }N=16,18,19,25,43,67,163,\\ \mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}&\text{ with }1\leq N\leq 9.\\ \end{cases}$
Moreover, for $\ell\in\{11,19,43,67,163\}$, the group $\mathbb{Z}/\ell\mathbb{Z}$ is realized only when $p=\ell$.
@@
By examining Table 2 in [3] and Table 4 in [19], we see that this occurs for only eleven different $j$-invariants. Using the LMFDB database, we analyze the field of definition of their $N$-torsion, where our elliptic curves correspond to the minimal quadratic twists of these $j$-invariants, or their $-3$-quadratic twists in some cases where this is more convenient (see Table 1).
In the case of $N=14$, there are two distinct $j$-invariants. When working with their minimal quadratic twists, the torsion growth observed on the LMFDB database shows that the field of definition for the $14$-torsion with an abelian Galois group is exactly $\mathbb{Q}(\zeta_{7})$. Let us focus on the unique $7$-isogeny. We see that its kernel is defined over $\mathbb{Q}(\zeta_{7})$. This means there exists a point $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ of order $7$, with $\mathbb{Q}(P)=\mathbb{Q}(\zeta_{7})$. Moreover, the set
@@
is the kernel of the unique (uniqueness follows from Theorem 4.2) $7$-isogeny (from $E$ to another elliptic curve), by Lemma 4.3. Considering the fact
@@
we see that $\mathbb{Q}(x_{0})$ has extension degree $3$ or $6$ over $\mathbb{Q}$. Therefore, $\mathbb{Q}(\zeta_{7})^{+}\subset\mathbb{Q}(x_{0})$ and $7$ is ramified at $\mathbb{Q}(x_{0})$.
Let $E_{d}(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/14\mathbb{Z}$ for some prime $p$ and square-free integer $d$. Then, $E_{d}$ has a $7$-isogeny by Lemma 4.3, and it is unique by Theorem 4.2. Thus, its kernel is generated by the point $P_{d}=(x_{0},y_{0}/\sqrt{d})\in E_{d}(\overline{\mathbb{Q}})_{\text{tors}}$, which is the image of $P$ on the quadratic twist $E_{d}$. Then, $P_{d}\in E_{d}(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ implies $\mathbb{Q}(x_{0})\subset\mathbb{Q}(\zeta_{p})$, implying that $7$ is ramified at $\mathbb{Q}(\zeta_{p})$. This implies that $14$-torsion is not realized over $\mathbb{Q}(\zeta_{p})$ when $p\neq 7$.
For the case of $N=15$, there are $4$ different $j$-invariants with a $15$-isogeny. Let $E/\mathbb{Q}$ have one of the corresponding $j$-invariants; then there is a point $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ of order $15$ such that the set
@@
is the kernel of a $15$-isogeny (from $E$ to some other elliptic curve). Moreover, this $15$-isogeny is unique due to Theorem 4.2. This means that $\langle P_{d}\rangle\subset E_{d}(\overline{\mathbb{Q}})_{\text{tors}}$ is the kernel of the unique $15$-isogeny of $E_{d}$. Here, as usual, $P_{d}=(x_{0},y_{0}/\sqrt{d})$ denotes the image of $P$ on the quadratic twist $E_{d}$. Then, to deal with all quadratic twists at once, we can pick one of them and consider the extension $\mathbb{Q}(x_{0})$, because it is the same extension for all quadratic twists of $E$. Our SageMath calculations show that $5$ is ramified at $\mathbb{Q}(x_{0})$ for all $4$ different $j$-invariants. Thus, $5$ must be ramified at $\mathbb{Q}(\zeta_{p})$, considering $\mathbb{Q}(x_{0})\subset\mathbb{Q}(\zeta_{p})$. This implies that $15$-torsion is not realized over $\mathbb{Q}(\zeta_{p})$ when $p\neq 5$.
For the case of $N=21$, there are $4$ different $j$-invariants with a $21$-isogeny. Let $E/\mathbb{Q}$ have one of the corresponding $j$-invariants; then there is a point $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ of order $21$ such that the set
@@
is the kernel of a $21$-isogeny (from $E$ to some other elliptic curve). Moreover, this $21$-isogeny is unique due to Theorem 4.2. This means that $\langle P_{d}\rangle\subset E_{d}(\overline{\mathbb{Q}})_{\text{tors}}$ is the kernel of the unique $21$-isogeny of $E_{d}$. Here, as usual, $P_{d}=(x_{0},y_{0}/\sqrt{d})$ denotes the image of $P$ on the quadratic twist $E_{d}$. Then, to deal with all quadratic twists at once, we can pick one of them and consider the extension $\mathbb{Q}(x_{0})$, because it is the same extension for all quadratic twists of $E$. Our SageMath calculations show that $3$ is ramified at $\mathbb{Q}(x_{0})$ for all $4$ different $j$-invariants. Thus, $3$ must be ramified at $\mathbb{Q}(\zeta_{p})$, considering $\mathbb{Q}(x_{0})\subset\mathbb{Q}(\zeta_{p})$. This implies that $21$-torsion is not realized over $\mathbb{Q}(\zeta_{p})$ when $p\neq 3$. By Theorem 2.2, we see that $E(\mathbb{Q}(\zeta_{3}))_{\text{tors}}\cong\mathbb{Z}/21\mathbb{Z}$ is not possible.
Finally, for the case of $N=27$, we note that there is only one rational $j$-invariant with a $27$-isogeny, namely when $j=-2^{15}\cdot 3\cdot 5^{3}$. This corresponds to the elliptic curve 27.a2 and its quadratic twists. Moreover, for any elliptic curve $E/\mathbb{Q}$ with $j(E)=-2^{15}\cdot 3\cdot 5^{3}$, the $27$-isogeny is unique due to Theorem 4.2. Let $E$ denote the minimal quadratic twist and $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ be a point of order $27$ generating the kernel of the $27$-isogeny.
@@
and $[\mathbb{Q}(\zeta_{27})^{+}:\mathbb{Q}]=9$, we can see that $\mathbb{Q}(x_{0})=\mathbb{Q}(\zeta_{27})^{+}$. In any quadratic twist $E_{d}$, the unique $27$-isogeny is generated by the point $P_{d}=(x,y/\sqrt{d})\in E_{d}(\overline{\mathbb{Q}})_{\text{tors}}$, hence the extension $\mathbb{Q}(x_{0})$ remains the same. Therefore, $27$-torsion is not realized over $\mathbb{Q}(\zeta_{p})$ for any prime $p$. ∎
We discuss the realization of torsion group structures other than those appearing in Mazur’s theorem in the next section.
With the theorem provided, we can classify the torsion groups of elliptic curves defined over $\mathbb{Q}(\zeta_{p})$ for primes $p>5$. To complete the classification, we can refer to the work of Gužvi’c and Vukorepa in [13], who classified the torsion subgroups for $p=5,7,11$. By incorporating their results, we can enhance the classification for these primes.
While our result offers a general classification for any $p>5$, the methods we used also provide a way to refine the list of realizable torsion subgroups for specific primes. By focusing on a particular prime $p$, one can further improve the classification and identify the torsion subgroups realized over $\mathbb{Q}(\zeta_{p})$. Although it is possible to obtain an exact list for each prime $p$, the task becomes increasingly difficult as $p$ grows larger.
@@
    /15,/16, $\mathbb{Z}/15\mathbb{Z},\quad\mathbb{Z}/16\mathbb{Z},\text{ and}\quad\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}.$
@@
    /13,/14,/18,/2×/14, $\mathbb{Z}/13\mathbb{Z},\quad\mathbb{Z}/14\mathbb{Z},\quad\mathbb{Z}/18\mathbb{Z},\quad\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/14\mathbb{Z},\text{ and}\quad\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/18\mathbb{Z}.$
@@
    /11,/25, $\mathbb{Z}/11\mathbb{Z},\quad\mathbb{Z}/25\mathbb{Z},\text{ and}\quad\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}.$
For the case of $p=3$, one can refer to Najman’s results in [22], where a classification of torsion subgroups over all quadratic fields is provided. This classification, along with the general methods discussed, helps refine the understanding of torsion subgroups for this particular prime.

### v6

@@ 4. Isogenies
  • If $C_{3}(E)=4$, then there exists a representative of the class of $E$ with a $\mathbb{Q}$-rational cyclic subgroup of order 272727, and $C_{2}(E)=1$ so $C(E)\leq 4$.
@@
This corollary is also used as a consequence of the Nйron–Ogg–Shafarevich criterion in the literature.
@@
Assume that $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/\ell\mathbb{Z}$ for some prime $p$ and some $\ell\in\{11,19,37,43,67,163\}$. By Lemma 4.3, $E$ must have an $\ell$-isogeny. In Chapter 7 of [3], the $j$-invariants of such elliptic curves are listed. In total, there are 777 different $j$-invariants, and we can also identify the Cremona labels of the corresponding elliptic curves. Using the LMFDB database, we observe that the conductors of these elliptic curves are equal to $N_{E}=\ell^{2}$ when working with the minimal quadratic twists of the corresponding $j$-invariants. This fact, combined with Corollary 4.6, implies that $\mathbb{Q}(E[\ell])$ is unramified at all primes other than $\ell$. For the remainder of the proof, the elliptic curve $E/\mathbb{Q}$ will always refer to the minimal quadratic twist of its $j$-invariant.
@@ 5. Relaxing the Conditions
In this section, we will discuss the implications of relaxing the conditions on our prime $p$ when classifying $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$. In our first theorem, we classified all possible torsion subgroups when $p-1$ is not divisible by 333, 444, or 555. Now, we will gradually relax these conditions. As a result, the number of groups appearing in the classification will increase. One of the first theorems we obtain is the following.
@@
Let $E/\mathbb{Q}$ be an elliptic curve. Let $p>3$ be a prime such that $p-1$ is not divisible by 333 or 444. Then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is either one of the groups from Mazur’s Theorem, or one of the following groups:
$\mathbb{Z}/11\mathbb{Z},\;\mathbb{Z}/16\mathbb{Z},\;\mathbb{Z}/25\mathbb{Z},\;\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z},\;\text{or }\;\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}.$
Among these, the groups $\mathbb{Z}/11\mathbb{Z}$, $\mathbb{Z}/25\mathbb{Z}$, $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$, and $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$ are realized for at least one prime $p$ satisfying the conditions given in the theorem. Moreover, the group $\mathbb{Z}/11\mathbb{Z}$ is realized only when $p=11$. The group $\mathbb{Z}/16\mathbb{Z}$ is theoretically possible under these assumptions, but no example is currently known.
@@
The proof is the same as that of Theorem 1.1. The only difference is in the application of Lemma 3.2 in the proof. Since we have removed the condition that $p-1$ is not divisible by 555, we cannot conclude that $\mathbb{Z}/11\mathbb{Z}$ or $\mathbb{Z}/25\mathbb{Z}$ are not realizable. We can use Lemma 4.7 to see that 111111-torsion is realized over $\mathbb{Q}(\zeta_{p})$ only when $p=11$ and only for 333 different $j$-invariants (see Table 2). The elliptic curve 11.a3 satisfies $E(\mathbb{Q}(\zeta_{11}))_{\text{tors}}\cong\mathbb{Z}/25\mathbb{Z}$. We could not find any other prime $p$ satisfying $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/25\mathbb{Z}$ for some elliptic curve $E/\mathbb{Q}$, and it is beyond our scope to show that no other primes exist. The groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$ and $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$ are realized for $p=59$ and $p=47$ respectively (see Section 6). ∎
@@
Let $E/\mathbb{Q}$ be an elliptic curve. Let $p>3$ be a prime such that $p-1$ is not divisible by 333. Then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is either one of the groups from Mazur’s Theorem, or one of the following groups:
$E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\simeq\begin{cases}\mathbb{Z}/N\mathbb{Z}&\text{ with }1\leq N\leq 13,\text{ or }N=15,16,25,\\ \mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}&\text{ with }1\leq N\leq 6,\text{ or }N=8,\\ \mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}.\\ \end{cases}$
Among these, all groups are realized for at least one prime $p$ satisfying the conditions of the theorem, with the exception of the group $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$, for which no example is currently known. Moreover, the group $\mathbb{Z}/11\mathbb{Z}$ is realized only when $p=11$. The groups $\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}$ and $\mathbb{Z}/15\mathbb{Z}$ are realized only when $p=5$.
@@
The proof is the same as in Theorem 1.1. The only difference lies in the use of Lemma 3.2 during the proof. Since we removed the condition that $p-1$ is not divisible by 444 or 555, we cannot rule out the possibility that $\mathbb{Z}/N\mathbb{Z}$ is realizable for $N=11,13,15,$ or 252525. Similarly, we cannot exclude the possibility of $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$. We also cannot rule out the cases $\mathbb{Z}/17\mathbb{Z}$ and $\mathbb{Z}/37\mathbb{Z}$ using Lemma 3.2, but Lemma 4.8 allows us to exclude them.
One final exception is that we cannot eliminate the $\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}$ case when $p=5$, because the Weil pairing argument does not apply. Indeed, the elliptic curve 11.a2 satisfies $E(\mathbb{Q}(\zeta_{5}))_{\text{tors}}\cong\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}$.
By Lemma 4.7, 111111-torsion is realized over $\mathbb{Q}(\zeta_{p})$ only when $p=11$. The elliptic curve 11.a3 satisfies $E(\mathbb{Q}(\zeta_{11}))_{\text{tors}}\cong\mathbb{Z}/25\mathbb{Z}$.
The elliptic curve 50.b3 satisfies $E(\mathbb{Q}(\zeta_{5}))_{\text{tors}}\cong\mathbb{Z}/15\mathbb{Z}$. In the proof of our next theorem, we show that $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/15\mathbb{Z}$ is realized only when $p=5$.
The elliptic curve 2890.a2 satisfies $E(\mathbb{Q}(\zeta_{17}))_{\text{tors}}\cong\mathbb{Z}/13\mathbb{Z}$. This data is available on the LMFDB. The field of definition of the 131313-torsion point is a quartic field with discriminant $17^{3}$ and Galois group isomorphic to $\mathbb{Z}/4\mathbb{Z}$. Using the ramified primes and the Kronecker–Weber theorem, we can deduce that this field is a subfield of $\mathbb{Q}(\zeta_{17})$. By Theorem 2.3, we conclude that the torsion subgroup over $\mathbb{Q}(\zeta_{17})$ must be isomorphic to $\mathbb{Z}/13\mathbb{Z}$, since a larger torsion subgroup is not possible.
We have no examples of $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$ for any prime $p$, and our methods are not effective in ruling out its existence when $p\equiv 1\pmod{4}$.
The groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$, $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$ and $\mathbb{Z}/16\mathbb{Z}$ are realized for $p=59$, $p=47$ and $p=41$ respectively (see Section 6). ∎
@@
Let $E/\mathbb{Q}$ be an elliptic curve. Let $p>3$ be a prime. Then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is either one of the groups from Mazur’s Theorem, or one of the following groups:
$E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\simeq\begin{cases}\mathbb{Z}/N\mathbb{Z}&\text{ with }1\leq N\leq 16,\text{ or }N=18,19,25,43,67,163,\\ \mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}&\text{ with }1\leq N\leq 9,\\ \mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}.\\ \end{cases}$
Among these, all groups are realized for at least one prime $p>3$, with the exception of the group $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$, for which no example is currently known. Moreover, for $\ell\in\{11,19,43,67,163\}$, the group $\mathbb{Z}/\ell\mathbb{Z}$ is realized only when $p=\ell$. The groups $\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}$ and $\mathbb{Z}/15\mathbb{Z}$ are realized only when $p=5$. The group $\mathbb{Z}/14\mathbb{Z}$ is realized only when $p=7$.
@@
By examining Table 2 in [3] and Table 4 in [19], we see that this occurs for only eleven different $j$-invariants. Using the LMFDB database, we analyze the field of definition of their $N$-torsion, where our elliptic curves correspond to the minimal quadratic twists of these $j$-invariants, or their 3-3- 3-quadratic twists in some cases where this is more convenient (see Table 2).
In the case of $N=14$, there are two distinct $j$-invariants. When working with their minimal quadratic twists, the torsion growth observed on the LMFDB database shows that the field of definition for the 141414-torsion with an abelian Galois group is exactly $\mathbb{Q}(\zeta_{7})$. Let us focus on the unique 777-isogeny. We see that its kernel is defined over $\mathbb{Q}(\zeta_{7})$. This means there exists a point $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ of order 777, with $\mathbb{Q}(P)=\mathbb{Q}(\zeta_{7})$. Moreover, the set
@@
is the kernel of the unique (uniqueness follows from Theorem 4.2) 777-isogeny (from $E$ to another elliptic curve), by Lemma 4.3. Considering the fact
@@
we see that $\mathbb{Q}(x_{0})$ has extension degree 333 or 666 over $\mathbb{Q}$. Therefore, $\mathbb{Q}(\zeta_{7})^{+}\subset\mathbb{Q}(x_{0})$ and 777 is ramified at $\mathbb{Q}(x_{0})$.
Let $E_{d}(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/14\mathbb{Z}$ for some prime $p$ and square-free integer $d$. Then, $E_{d}$ has a 777-isogeny by Lemma 4.3, and it is unique by Theorem 4.2. Thus, its kernel is generated by the point $P_{d}=(x_{0},y_{0}/\sqrt{d})\in E_{d}(\overline{\mathbb{Q}})_{\text{tors}}$, which is the image of $P$ on the quadratic twist $E_{d}$. Then, $P_{d}\in E_{d}(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ implies $\mathbb{Q}(x_{0})\subset\mathbb{Q}(\zeta_{p})$, implying that 777 is ramified at $\mathbb{Q}(\zeta_{p})$. This implies that 141414-torsion is not realized over $\mathbb{Q}(\zeta_{p})$ when $p\neq 7$.
For the case of $N=15$, there are 444 different $j$-invariants with a 151515-isogeny. Let $E/\mathbb{Q}$ have one of the corresponding $j$-invariants; then there is a point $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ of order 151515 such that the set
@@
is the kernel of a 151515-isogeny (from $E$ to some other elliptic curve). Moreover, this 151515-isogeny is unique due to Theorem 4.2. This means that $\langle P_{d}\rangle\subset E_{d}(\overline{\mathbb{Q}})_{\text{tors}}$ is the kernel of the unique 151515-isogeny of $E_{d}$. Here, as usual, $P_{d}=(x_{0},y_{0}/\sqrt{d})$ denotes the image of $P$ on the quadratic twist $E_{d}$. Then, to deal with all quadratic twists at once, we can pick one of them and consider the extension $\mathbb{Q}(x_{0})$, because it is the same extension for all quadratic twists of $E$. Our SageMath calculations show that 555 is ramified at $\mathbb{Q}(x_{0})$ for all 444 different $j$-invariants. Thus, 555 must be ramified at $\mathbb{Q}(\zeta_{p})$, considering $\mathbb{Q}(x_{0})\subset\mathbb{Q}(\zeta_{p})$. This implies that 151515-torsion is not realized over $\mathbb{Q}(\zeta_{p})$ when $p\neq 5$.
For the case of $N=21$, there are 444 different $j$-invariants with a 212121-isogeny. Let $E/\mathbb{Q}$ have one of the corresponding $j$-invariants; then there is a point $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ of order 212121 such that the set
@@
is the kernel of a 212121-isogeny (from $E$ to some other elliptic curve). Moreover, this 212121-isogeny is unique due to Theorem 4.2. This means that $\langle P_{d}\rangle\subset E_{d}(\overline{\mathbb{Q}})_{\text{tors}}$ is the kernel of the unique 212121-isogeny of $E_{d}$. Here, as usual, $P_{d}=(x_{0},y_{0}/\sqrt{d})$ denotes the image of $P$ on the quadratic twist $E_{d}$. Then, to deal with all quadratic twists at once, we can pick one of them and consider the extension $\mathbb{Q}(x_{0})$, because it is the same extension for all quadratic twists of $E$. Our SageMath calculations show that 333 is ramified at $\mathbb{Q}(x_{0})$ for all 444 different $j$-invariants. Thus, 333 must be ramified at $\mathbb{Q}(\zeta_{p})$, considering $\mathbb{Q}(x_{0})\subset\mathbb{Q}(\zeta_{p})$. This implies that 212121-torsion is not realized over $\mathbb{Q}(\zeta_{p})$ when $p\neq 3$. By Theorem 2.2, we see that $E(\mathbb{Q}(\zeta_{3}))_{\text{tors}}\cong\mathbb{Z}/21\mathbb{Z}$ is not possible.
Finally, for the case of $N=27$, we note that there is only one rational $j$-invariant with a 272727-isogeny, namely when $j=-2^{15}\cdot 3\cdot 5^{3}$. This corresponds to the elliptic curve 27.a2 and its quadratic twists. Moreover, for any elliptic curve $E/\mathbb{Q}$ with $j(E)=-2^{15}\cdot 3\cdot 5^{3}$, the 272727-isogeny is unique due to Theorem 4.2. Let $E$ denote the minimal quadratic twist and $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ be a point of order 272727 generating the kernel of the 272727-isogeny.
@@
and $[\mathbb{Q}(\zeta_{27})^{+}:\mathbb{Q}]=9$, we can see that $\mathbb{Q}(x_{0})=\mathbb{Q}(\zeta_{27})^{+}$. In any quadratic twist $E_{d}$, the unique 272727-isogeny is generated by the point $P_{d}=(x,y/\sqrt{d})\in E_{d}(\overline{\mathbb{Q}})_{\text{tors}}$, hence the extension $\mathbb{Q}(x_{0})$ remains the same. Therefore, 272727-torsion is not realized over $\mathbb{Q}(\zeta_{p})$ for any prime $p$.
The groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$, $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$ and $\mathbb{Z}/16\mathbb{Z}$ are realized for $p=59$, $p=47$ and $p=41$ respectively (see Section 6). ∎
We discuss, in more detail, the realization of torsion group structures other than those appearing in Mazur’s theorem in the next section.
With the theorem provided, we can classify the torsion subgroups of elliptic curves defined over $\mathbb{Q}(\zeta_{p})$ for primes $p>3$. To complete the classification, we can refer to the work of Gužvić and Vukorepa in [13], who classified the torsion subgroups for $p=5,7,11$. By incorporating their results, we can enhance the classification for these primes.
While our result offers a general classification for any $p>3$, the methods we used also provide a way to refine the list of realizable torsion subgroups for specific primes. By focusing on a particular prime $p$, one can further improve the classification and identify the torsion subgroups realized over $\mathbb{Q}(\zeta_{p})$. Although it is possible to obtain an exact list for each prime $p$, the task becomes increasingly difficult as $p$ grows larger.
@@
    $\mathbb{Z}/15\mathbb{Z},\quad\mathbb{Z}/16\mathbb{Z},\text{ and}\quad\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}.$
@@
    $\mathbb{Z}/13\mathbb{Z},\quad\mathbb{Z}/14\mathbb{Z},\quad\mathbb{Z}/18\mathbb{Z},\quad\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/14\mathbb{Z},\text{ and}\quad\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/18\mathbb{Z}.$
@@
    $\mathbb{Z}/11\mathbb{Z},\quad\mathbb{Z}/25\mathbb{Z},\text{ and}\quad\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}.$
For the case of $p=3$, one can refer to Najman’s results in [23], where he provides a general classification of torsion subgroups over all quadratic fields. In a separate work [22], he gives a more specific classification for the field $\mathbb{Q}(\sqrt{-3})$ (and also for $\mathbb{Q}(\sqrt{-1})$), which directly addresses the case $p=3$. This classification, along with the general methods discussed, helps refine the understanding of torsion subgroups for this particular prime.